# Thai Large Language Models

### v1

@@ 5.1 End-to-End Model Architecture
The architecture of Typhoon2-Audio, illustrated in Figure 7, processes both text and audio/speech121212In this section, the terms “audio” and “speech” will be used interchangeably. The input may consist of human speech or general audio events, but the output is speech generated by a single voice. modalities, integrating pre-trained components for text and speech handling. It starts with an encoding module comprising a text tokenizer for converting textual input into representations and a speech encoder that embeds speech and audio-event inputs into a shared representation space. These embeddings are passed to an LLM for reasoning and generating outputs. The speech generation module then converts LLM outputs into: (1) text via a language model head, and (2) speech via a speech decoder and unit vocoder, producing speech output. Text and speech outputs can be decoded in parallel. The encoding module adopts the design of Typhoon-Audio (Manakul et al., 2024) based on SALMONN (Tang et al., 2024a), while the speech generation module follows Llama-Omni (Fang et al., 2024).
@@ 5.2.1 Speech Encoder
Encoder Architecture: Our model leverages the SALMONN architecture (Tang et al., 2024a), employing BEATs for audio event understanding, Whisper Encoder for speech understanding, and Q-Former with an MLP to align speech and text representations With Thai and English as target languages, our model is based on an LLM from the Typhoon series (e.g., Typhoon 1.5, Typhoon 2), Whisper-large-v3 fine-tuned to Thai (Aung et al., 2024) coupled with BEATs (Chen et al., 2023) as the audio encoder, and Q-Former (Li et al., 2023a) trained from scratch as the adapter. Note that we also examine other variants of LLM and audio encoder backbones in Section 5.2.
@@ 5.2.4 Results and Findings
Remarks: The majority of experiments in Section 5.2.4 were conducted prior to the development of the Typhoon2 Text series. Thus, the best performing 8B Typhoon backbone at the time, Typhoon1.5 "llama-3-typhoon-v1.5-8b-instruct", was selected as the main LLM backbone. In the final experiment (discussed in Finding 5), we provide the performance of the new audio model with Typhoon2 "typhoon-2-llama-31-8b-instruct-beta-v1".
@@ 5.3.4 Data
Mix-1: This data mixture includes 28.7K examples derived from Thai self-instruct (8.7K) and a translated Alpaca subset (20K). Responses are regenerated using GPT-4o-mini in a conversational style, following Llama-Omni’s setup.131313It was intended to fine-tune the speech encoder + LLM for more conversational responses, but experiments showed it degraded text generation, so the original speech encoder + LLM was retained.
Mix-2: This mixture comprises 220K examples derived from: (1) 75K re-written Thai SFT data,141414Dataset sourced from https://huggingface.co/datasets/Suraponn/thai_instruction_sft. (2) responses generated from re-written instructions, and (3) 70K examples of sentences using Thai unique names (e.g., companies, locations).151515Thai company names were limited, so this portion was upsampled. Instructions and responses follow the same GPT-4o-mini conversational generation style as Mix-1 but on a larger scale.
@@ 6 Conclusions
This technical report introduced Typhoon 2, a series of Thai large language models, comprising text models in multiple sizes (both base and instruction-tuned variants) and multimodal models for vision and audio tasks. Our evaluation demonstrates superior performance across a majority of evaluated tasks, including math and reasoning. Typhoon2-Text models have an extended context length of up to 128,000 tokens, compared to 8,192 tokens of Typhoon 1.5. Typhoon2-Text also features function calling capabilities, achieving state-of-the-art results in our experiments. Additionally, we developed a safety classifier that delivers state-of-the-art performance specifically for Thai.
@@ 7 Acknowledgments
We would like to thank Thanapong Boontaeng, Mukaya Panich, and other members of SCB 10X for their support and guidance on the project. We also extend our appreciation to Kaweewut Temphuwapat and SCBX Innovation Team for their support, resources, and valuable insights in advancing this AI development. Lastly, we thank the global and local AI communities for open-sourcing invaluable resources and sharing knowledge.

### v2

@@ 5.1 End-to-End Model Architecture
The architecture of Typhoon2-Audio, illustrated in Figure 7, processes both text and audio/speech131313In this section, the terms “audio” and “speech” will be used interchangeably. The input may consist of human speech or general audio events, but the output is speech generated by a single voice. modalities, integrating pre-trained components for text and speech handling. It starts with an encoding module comprising a text tokenizer for converting textual input into representations and a speech encoder that embeds speech and audio-event inputs into a shared representation space. These embeddings are passed to an LLM for reasoning and generating outputs. The speech generation module then converts LLM outputs into: (1) text via a language model head, and (2) speech via a speech decoder and unit vocoder, producing speech output. Text and speech outputs can be decoded in parallel. The encoding module adopts the design of Typhoon-Audio (Manakul et al., 2024) based on SALMONN (Tang et al., 2024a), while the speech generation module follows Llama-Omni (Fang et al., 2024).
@@ 5.2.1 Speech Encoder
Encoder Architecture: Our model leverages the SALMONN architecture (Tang et al., 2024a), employing Whisper Encoder for speech understanding, BEATs for audio event understanding, and Q-Former with an MLP to align speech and text representations With Thai and English as target languages, our model is based on an LLM from the Typhoon series (e.g., Typhoon 1.5, Typhoon 2), Whisper-large-v3 fine-tuned to Thai (Aung et al., 2024) coupled with BEATs (Chen et al., 2023) as the audio encoder, and Q-Former (Li et al., 2023a) trained from scratch as the adapter. Note that we also examine other variants of LLM and audio encoder backbones in Section 5.2.
@@ 5.2.4 Results and Findings
Remarks: The majority of experiments in Section 5.2.4 were conducted prior to the development of the Typhoon2 Text series. Thus, the best performing 8B Typhoon backbone at the time, Typhoon1.5 "llama-3-typhoon-v1.5-8b-instruct", was selected as the main LLM backbone. In the final experiment (discussed in Finding 4), we provide the performance of the new audio model with Typhoon2 "typhoon-2-llama-31-8b-instruct-beta-v1".
@@ 5.3.4 Data
Mix-1: This data mixture includes 28.7K examples derived from Thai self-instruct (8.7K) and a translated Alpaca subset (20K). Responses are regenerated using GPT-4o-mini in a conversational style, following Llama-Omni’s setup.141414It was intended to fine-tune the speech encoder + LLM for more conversational responses, but experiments showed it degraded text generation, so the original speech encoder + LLM was retained.
Mix-2: This mixture comprises 220K examples derived from: (1) 75K re-written Thai SFT data,151515Dataset sourced from https://huggingface.co/datasets/Suraponn/thai_instruction_sft. (2) responses generated from re-written instructions, and (3) 70K examples of sentences using Thai unique names (e.g., companies, locations).161616Thai company names were limited, so this portion was upsampled. Instructions and responses follow the same GPT-4o-mini conversational generation style as Mix-1 but on a larger scale.
@@ 6 Conclusions
This technical report introduced Typhoon 2, a series of Thai LLMs, comprising text models in multiple sizes (both base and instruction-tuned variants) and multimodal models for vision and audio tasks. Our evaluation demonstrates superior performance across a majority of evaluated tasks, including math and reasoning. Typhoon2-Text models have an extended context length of up to 100,000 tokens, compared to 8,192 tokens of Typhoon 1.5. Typhoon2-Text also features function calling capabilities, achieving state-of-the-art results. Additionally, we include a safety classifier that delivers state-of-the-art performance specifically for Thai.
@@ 7 Acknowledgments
We thank Thanapong Boontaeng for his insightful advice and feedback on this technical report and the overall project. We also thank Mukaya Panich and other members of SCB 10X for their invaluable support and guidance throughout the project. We also extend our appreciation to Kaweewut Temphuwapat and the SCBX Innovation Team for their support, resources, and valuable insights that have made this project possible. Lastly, we are grateful to the global and local AI communities for open-sourcing resources and knowledge.